Methanol (CH3OH) burns in air according to the equation: 2〖CH〗_3OH(l) + 3O_3 (g) 4H_2 0 (l) + 2〖CO〗_2 (g) ΔH = -1452.8 kJ/mol

How much heat is produced by the complete reaction of 1.26 kg CH3OH?

Molar mass CH3OH = 32.05 g/mol
1
Expert's answer
2020-10-20T08:52:06-0400

n (CH3OH) = m/M = 1 260 /32.05 = 39.3 mol

ΔH = -1 452.8 kJ/mol x 39.3 mol = - 57 115 kJ